# Follow-Bench: A Unified Motion Planning Benchmark for Socially-Aware Robot Person Following

### v2

@@ II-C RPF-oriented Motion Planning
RPF-oriented motion planning must simultaneously account for safety and comfort. It generally follows the standard navigation pipeline—comprising global path planning and local motion planning—while incorporating additional human-aware functions or constraints. A representative framework is illustrated in Fig. 2, which includes both global and local planners. Global planners typically rely on a prior map [48, 52, 46, 61, 32, 66] to estimate the target’s long-term destination and to compute a global path. These methods often emphasize safety, particularly in re-finding the target after occlusion, in order to maintain continuous visibility [52, 66]. However, constructing a reliable prior map is impractical in many real-world scenarios, especially in large-scale or dynamic environments. Given the dynamic nature of RPF, most approaches therefore focus on local planners, which are responsible for generating safe and socially comfortable person-following behaviors. For this reason, our review primarily concentrates on local planning methods. A summary of representative approaches and their attributes is provided in Table III. For simplicity, we hereafter refer to RPF-oriented local motion planners as RPF planners.
@@
Current RPF planners typically navigate to a goal based on the target’s most recent position, which has been demonstrated more human-like and socially compliant than replicating the target’s historical trajectory [67]. The most straightforward strategy is to predefine a relative goal with respect to the target and use a PID controller to reach it [27, 5, 6, 68]. Some methods [27, 32] further incorporate proxemic constraints to determine socially appropriate accompanying positions. Such constraints generally include distance and relative position. The distance constraint [27] is usually defined as a comfortable following distance within the social space (Fig. 1), while position preferences are commonly categorized into frontal [29], side-by-side [32], and back following [27]. Among these, back following is the most prevalent due to its unobtrusive nature. Frontal following is less common since it requires accurate prediction of the target’s future movements while avoiding obstacles, whereas side-by-side following is typically adopted in socially interactive scenarios.
@@
Beyond proxemic constraints, collision avoidance is essential. To address this, more sophisticated local planning techniques have been proposed, including dynamic window approach (DWA)[69, 31, 64], SFM[14, 70], model predictive control (MPC)[24, 7, 33, 71, 72], and reinforcement learning (RL)[73, 47, 29, 44]. For example, Dang et al.[64] enhance the DWA by incorporating the desired person-following direction into the velocity sampling process, along with collision constraints and smoothness considerations. SFM methods model the robot’s movement using attractive forces toward the target (or their destination), repulsive forces to avoid obstacles, and social forces to respect interpersonal distance from pedestrians[70].
@@
In addition to proxemics compliance and collision avoidance, maintaining target visibility is another critical aspect of RPF safety. Existing methods address this through active planning or person search. Active planning anticipates occlusion events and proactively adjusts the robot’s position to maintain line-of-sight. For example, Yao et al.[26] address corner occlusion by predicting the target’s turn and switching from back-following to side-following via a PD controller. While effective in specific cases, this method lacks flexibility to diverse occlusion scenarios. Wang et al.[65] propose a more flexible MPC-based approach that introduces human-human and human-obstacle occlusion constraints. However, while maximizing visibility, it may compromise proxemic comfort, resulting in overly distant following behaviors.
@@ III-A Overview
To comprehensively evaluate the local motion planning capabilities of various RPF planners within a unified framework, we present Follow-Bench. This benchmark extends a Python-based 2D navigation simulator[80] by incorporating realistic interaction simulations, as well as human visibility checks. Several features are inspired by existing social navigation simulators[17, 81]. Here, we adopt a lightweight simulator instead of high-fidelity simulators for its simplicity and compatibility with CPU-only execution. Moreover, it is worth noting that the interaction mechanisms in many high-fidelity simulators[56, 18] are also fundamentally 2D-based: they first generate socially-aware 2D human trajectories and subsequently augment them with 3D animations for enhanced visualization.
Follow-Bench offers diverse scenarios (Sec. III-B), including various target trajectory patterns, dynamic-crowd flow, and environmental layouts, where the robot is tasked with following a specific individual while avoiding obstacles and ensuring safety and comfort for both the target and surrounding people. It includes an exhaustively re-implemented RPF planners (Sec. III-C), along with evaluation metrics (Sec. III-D) that emphasize safety and comfort. The benchmark is designed to be highly extensible, allowing for the addition of new maps, obstacles, agents and agents’ trajectories and motion attributes.
@@ III-B Scenarios
In this benchmark, we construct a diverse set of scenarios spanning variations in target trajectories, environmental layouts, and dynamic crowd patterns, drawing upon representative RPF settings summarized in Table I.
@@
Target trajectory patterns include two-triangle, two-square, U-turn, figure-eight (‘8’) patterns, L-turns with varying angles (30∘, 45∘, and 60∘), and back-and-forth walking (Fig. 3).
Dynamic-Crowd flows cover parallel crossing, perpendicular crossing, circular crossing, and randomly distributed pedestrian flows (Fig. 4).
Environmental layouts comprise corridors, doorways, intersections, and cluttered spaces populated with dynamic pedestrians (Fig. 5).
These scenario types are widely adopted in RPF-related studies for evaluating the adaptive intelligence of planners. Building on them, we also assess different preferred following positions, including back-following [53] and side-following [46] (Fig. 3).
In all scenarios, human trajectories are generated via an A* global planner over predefined waypoints, while avoiding obstacles based on ORCA or SFM. The baseline configuration models each human as an omni-directional dynamic cylinder (radius: , max speed: ) and the robot as a differential-drive platform () equipped with a 2D LiDAR ( range), subject to translational velocity and acceleration limits. The framework is inherently modular: agent geometry, sensing profile, and motion dynamics can be readily reconfigured to support alternative embodiments, including omni-directional, differential-drive, and Ackermann steering systems.
The evaluated scenarios are summarized in Table IV. To quantify the complexity of socially aware scenarios, we adopt the characterization framework of Stratton et al. [16], evaluating planner performance under varying human densities and environmental occupancy conditions (i.e., passageway widths or obstacle numbers). For dynamic-crowd flows, we define six number levels—5, 10, 15, 20, 25, and 30. For environmental layout scenarios, number levels are set individually according to spatial constraints: Corridor (4, 12, 20, 28) with a width; Intersection (2, 10, 18, 26) with a width; Doorway (2, 6, 10, 14) with a doorway width; and Cluttered space (10, 20, 30, 40) with 30 random obstacles. Furthermore, we vary occupancy conditions by either narrowing passageway widths or increasing obstacle numbers: Corridor (6.8, 6.2, 5.6, 5.0) with 20 humans, Intersection (6.0, 5.4, 4.8, 4.2) with 18 humans, Doorway (3.4, 3.1, 2.8, 2.5) with 10 humans, and Cluttered space (40, 30, 20, 10 obstacles) with 30 humans, respectively. The maximum pedestrian densities, obstacle numbers, and minimum passage widths are determined empirically by observing conditions under which simulated pedestrians become prone to congestion.
In addition, we also evaluate the impact of different person-following positions (back-following and side-following, and different following distances—1.0, 1.5, 2.0, 2.5 meters) on planner performance. These variations aim to assess planners’ adaptability to different social preferences and proxemic requirements. Specifically, we evaluate back- and side-following with following distance in the above all scenarios. While under the variation of following distance, the human number of dynamic-crowd flows is set to 20, and environmental layouts are set to Corridor (20 humans, width), Intersection (18 humans, width), Doorway (10 humans, width), and Cluttered space (30 humans, 30 obstacles).
The benchmark is continuously expanded with new scenarios, and the scenario-generation scripts are publicly available, enabling researchers to freely adjust parameters such as human count, corridor width, and overall scenario scale.
III-C Baselines
@@
As reviewed in Sec. II-C and summarized in Table III, numerous RPF planners have been proposed. However, most implementations are not publicly available and, critically, few address obstacle avoidance and proxemics in an integrated manner. To this end, we re-implement a representative set of well-studied RPF planners validated in real-world dynamic environments and explicitly designed to satisfy both safety and comfort requirements. The re-implemented RPF planners include a SFM-based planner [61], a MPC-based planner [80], a DWA-based planner [64], and an enhanced DWA-based planner that incorporates target-trajectory prediction (DWA w/ Traj.).
@@
Proxemic constraints are considered standard practice in recent RPF literature [27, 33, 7]. Accordingly, each planner maintains a “comfortable” following distance of 1.5 m within the target’s personal space. To mitigate target obstruction, we adopt adaptive following-position adjustments similar to [7, 75]. In addition, each baseline is equipped with a standard RPF search module [77, 82, 83] that re-finds the target after loss by navigating to the Kalman-filter-predicted position. Given the prominence of MPC-based approaches, we additionally reproduce two MPC variants: (i) MPC w/ trajectory prediction (MPC w/ Traj.) [7], which guides the robot along a predicted target trajectory inferred from historical motion. This planner halts when the predicted trajectory exhibits high uncertainty, thereby prioritizing safety; and (ii) MPC with dynamic search fields (MPC w/ DS.) [53], which employs a dynamic search field to intelligently re-find the target after occlusion—either by overtaking occluders or by temporarily following them until an overtaking opportunity arises.
@@
The current release of Follow-Bench focuses on these map-independent local planners, evaluating their fundamental ability to avoid collisions and stably follow the target across diverse environments. To decouple motion-planning evaluation from perception, we assume perfect knowledge of the target and pedestrian positions whenever visible, with occlusion handled by the simulator’s detection logic. Static obstacles are detected using a clustering-based method applied to 2D LiDAR input, following [80]. For completeness, each baseline is implemented in both differential-drive and omnidirectional versions.
@@ III-D Evaluation Protocol and Metrics
Using the re-implemented RPF planners described above, we evaluate their performance across the benchmark scenarios introduced in Sec. III-B. The simulation timestep is fixed at 0.1 s, meaning each step corresponds to 0.1 seconds of simulated time. For each scenario, we first determine the minimum number of simulation steps, , required for the target person to complete their trajectory without the robot present, while accounting for interactions with other pedestrians. Each scenario is then simulated with the robot included for a total of steps—a duration that ensures the target has ample time to finish the trajectory.
Each scenario is tested with 100 randomized trials, with surrounding pedestrians’ trajectories and any static obstacles varied across trials. The robot is always initialized at the preferred following position relative to the target. For each trial, we compute the evaluation metrics listed in Table II. Specifically, search success is defined as the robot successfully re-finding the target within steps after each target loss until the trial ends. A successful trial is defined as one that completes without any collisions and in which the robot recovers tracking within steps of the target’s completion.
@@ IV Analysis of RPF planners
Based on the proposed benchmark, we conduct a series of experiments to evaluate the performance of re-implemented RPF planners. The experiments are designed to analyze how these planners perform under all scenarios (Sec. IV-A), different person-following distances and positions, human densities and environmental occupancy conditions (Sec. IV-B). Here, we simulate human-object and human-human interaction based on ORCA. For simplicity, we show part of the evaluation metrics here, detailed metric evaluation can be found in the released code. In addition, Appendix includes complete evaluation results of all experiments in each scenario.
@@ V-B1 More Context-rich Environments
Currently, our Follow-Bench encompasses the most representative scenarios in RPF research, including various target trajectory patterns, dynamic-crowd flows, and environmental layouts. Nevertheless, it primarily evaluates basic motion-planning capabilities in terms of safety and comfort. Future extensions could introduce more context-rich environments, such as hospitals, factories, museums, households, and offices. In these settings, the target person may exhibit application-specific behaviors and interactions with the environment (e.g., visiting a patient, collaborating with colleagues, or shopping), while the robot could be assigned complementary task-oriented goals during following (e.g., delivering medicine, carrying luggage, or guiding to a destination). Such scenarios would demand more sophisticated reasoning and decision-making from RPF planners.
In addition, the current benchmark is implemented in a 2D simulator for ease of evaluating and developing RPF planners. A high-fidelity and more realistic 3D simulator (e.g., IsaacSim [97]) would enable a richer evaluation of RPF systems under perceptual challenges, including humans’ trajectory prediction, visual occlusion [53], visual localization from quadruped platforms [41], and semantic understanding [36].
@@ VI Conclusions
Robot person following is a fundamental capability in human-robot interaction, with applications spanning personal assistance, security, service, and industrial robotics. This paper provides a comprehensive and structured survey of existing RPF research, with a particular emphasis on two widely recognized and critical criteria: safety and comfort. Guided by these objectives, we review representative scenarios, evaluation metrics, and motion planners. To enable systematic benchmarking of reviewed planners, we introduce Follow-Bench, a unified benchmark implemented in a lightweight simulator that models socially aware interactions, including diverse target trajectory patterns, dynamic-crowd flows, and environmental layouts. To the best of our knowledge, this is the first benchmark to explicitly highlight two essential requirements for RPF: (i) continuously and naturally following a target person, and (ii) proactively avoiding obstacles and pedestrians in socially interactive environments. Follow-Bench allows researchers to develop and evaluate new RPF planners, with flexible support for adding maps, obstacles, agents, metrics, and following strategies. All code and deployment scripts are publicly released.

### v3

@@ II-C RPF-oriented Motion Planning
RPF-oriented motion planning must simultaneously account for safety and comfort. It generally follows the standard navigation pipeline—comprising global path planning and local motion planning—while incorporating additional human-aware functions or constraints. A representative framework is illustrated in Fig. 2, which includes both global and local planners. Global planners typically rely on a prior map [45, 50, 43, 60, 29, 65] to estimate the target’s long-term destination and to compute a global path. These methods often emphasize safety, particularly in re-finding the target after occlusion, in order to maintain continuous visibility [50, 65]. However, constructing a reliable prior map is impractical in many real-world scenarios, especially in large-scale or dynamic environments.
Given the dynamic nature of RPF, most approaches therefore focus on local planners, which are responsible for generating safe and socially compliant person-following behaviors. For this reason, our review primarily concentrates on local planning methods. In contrast to previous surveys [1, 4] that mainly discuss collision avoidance and proxemic considerations, we further analyze how recent methods incorporate active occlusion avoidance, recovery after occlusion, and adaptive following position to enhance both safety and comfort. A summary of representative approaches and their attributes is presented in Table III. For simplicity, we hereafter refer to RPF-oriented local motion planners as RPF planners.
@@
Current RPF planners typically navigate to a goal based on the target’s most recent position, which has been demonstrated more human-like and socially compliant than replicating the target’s historical trajectory [66]. The most straightforward strategy is to predefine a relative position with respect to the target and use a PID controller to reach and maintain it [24, 5, 6, 67]. For comfort consideration, some methods [24, 29] further incorporate proxemic constraints to regulate the robot’s accompanying position. These constraints generally specify the desired distance and angle between the robot and the target person. The distance constraint [24] is usually defined as a comfortable following distance within the social space (Fig. 1), while angular preferences are commonly categorized into frontal [26], side-by-side [29], and back following [24]. Among these, back following is the most prevalent due to its unobtrusive nature. Frontal following is less common since it requires accurate prediction of the target’s future movements while avoiding obstacles, whereas side-by-side following is typically adopted in socially interactive scenarios.
@@
Beyond the proxemic constraints considered for comfort, collision avoidance remains a fundamental safety requirement. To simultaneously satisfy both proxemic and collision-avoidance constraints, researchers have developed increasingly sophisticated local planning techniques, including dynamic window approach (DWA)[68, 28, 63], SFM[14, 69], model predictive control (MPC)[21, 7, 30, 70, 71], and reinforcement learning (RL)[72, 46, 26, 41]. For example, Dang et al.[63] enhance the DWA by incorporating the desired person-following direction into the velocity sampling process, along with collision constraints and smoothness considerations. SFM methods model the robot’s movement using attractive forces toward the target (or their destination), repulsive forces to avoid obstacles, and social forces to respect interpersonal distance from pedestrians[69].
@@
Beyond collision avoidance, occlusion avoidance represents another critical dimension of RPF safety. Occlusions caused by static structures (e.g., walls or furniture) or dynamic obstacles (e.g., surrounding pedestrians) may lead to temporary loss of the target, resulting in discontinuous or uncomfortable following behaviors. Existing approaches typically address this challenge via either active planning[75, 23, 64] or person search[48, 76, 77, 78, 79] strategies.
Active planning anticipates occlusion events and proactively adjusts the robot’s position to maintain line-of-sight. For example, Yao et al.[23] address corner occlusion by predicting the target’s turn and switching from back-following to side-following via a PD controller. While effective in specific cases, this method lacks flexibility to diverse occlusion scenarios. Wang et al.[64] propose a more flexible MPC-based approach that introduces human-human and human-obstacle occlusion constraints. However, while maximizing visibility, it may compromise proxemic comfort, resulting in overly distant following behaviors.
@@ III-A Overview
To comprehensively evaluate the local motion planning capabilities of various RPF planners within a unified framework, we present Follow-Bench. This benchmark extends a Python-based 2D navigation simulator[80] by incorporating realistic interaction simulations, as well as human visibility checks. Several features are inspired by existing social navigation simulators[17, 81].
Follow-Bench provides a diverse set of scenarios (Sec. III-B), encompassing various target trajectory patterns, crowd dynamics, and environmental layouts. In each scenario, the robot is required to follow a designated individual while avoiding obstacles and ensuring both safety and comfort for the target and surrounding pedestrians. These scenarios are selected from the reviewed studies (Sec. II-A), representing typical yet challenging situations commonly encountered in RPF tasks. Furthermore, Follow-Bench includes a comprehensive re-implementation of representative RPF planners (Sec. III-C) and a set of evaluation metrics (Sec. III-D) emphasizing safety and comfort. The benchmark is designed with high extensibility, enabling the seamless integration of new maps, obstacles, agents, as well as agent trajectories and motion attributes.
@@ III-B Scenarios
Follow-Bench encompasses a diverse suite of 16 scenario types, systematically designed to capture variations in target trajectory patterns, crowd dynamics, and environmental layouts—drawing upon representative RPF scenarios summarized in Table I.
@@
Target trajectory patterns include two-triangle[7, 30], two-square[7, 20, 21, 22, 23, 24], U-turn[7, 25, 26, 27, 28, 29], figure-eight (‘8’) patterns[7, 25], L-turns with varying angles (30∘, 45∘, and 60∘)[7], and back-and-forth walking[7]. Representative examples are shown in Fig. 3. These scenarios involve abundant turning behaviors, designed to independently evaluate an RPF planner’s agility, trajectory smoothness, and adaptability to target motion changes.
Crowd dynamics cover parallel crossing[35, 36, 42, 43], perpendicular crossing[35, 42, 43], circular crossing[44], and random crowds[13, 14]. Due to human-human interactions, pedestrian motions in these scenarios are highly dynamic and uncertain, often resulting in potential collisions and occlusions. They are therefore employed to evaluate the planner’s capability in collision avoidance, occlusion recovery, and maintaining socially appropriate interpersonal distances in human-populated environments. Representative examples are shown in Fig. 4.
Environmental layouts comprise corridors[21, 28, 24, 35, 36], doorways[13, 37], intersections[34, 36], and cluttered spaces populated with dynamic pedestrians[20, 37]. These scenarios combine both dynamic and static obstacles, posing higher-level challenges where the robot must traverse constrained environments while ensuring collision-free motion and socially comfortable distances from both the target and nearby pedestrians. Representative cases are depicted in Fig. 5.
Building upon the aforementioned scenarios, Follow-Bench further supports evaluation across different following angles, including back-following [48] and side-following [43], as illustrated in Fig. 3. To systematically quantify the complexity of RPF tasks, we adopt the characterization framework proposed by Stratton et al. [16], evaluating planner performance under varying numbers of humans and environmental occupancy levels (e.g., passageway widths and obstacle counts). Moreover, we investigate the impact of different following distances to analyze each planner’s adaptability to diverse social preferences and proxemic constraints. These configuration variants within each scenario enable systematic assessment of a planner’s adaptability to varying levels of complexity—including collision and occlusion risks, as well as the maintenance of socially appropriate interpersonal distances. The detailed configuration specifications are summarized in Table IV.
In all scenarios, target and pedestrian trajectories are generated using an A* global planner over predefined waypoints, while local interactions and obstacle avoidance are governed by either ORCA or SFM. Further implementation details of the simulation environment are provided in the Appendix.
III-C RPF planners
@@
We re-implement a representative set of well-studied RPF planners validated in real-world dynamic environments and explicitly designed to satisfy both safety and comfort requirements. The re-implemented RPF planners include a SFM-based planner [60], a MPC-based planner [80], a DWA-based planner [63], and an enhanced DWA-based planner that incorporates target-trajectory prediction (DWA w/ Traj.).
@@
Proxemic constraints are considered standard practice in recent RPF literature [24, 30, 7]. Accordingly, each planner maintains a “comfortable” following distance of 1.5 m within the target’s personal space. To mitigate target obstruction, we adopt adaptive following-position adjustments similar to [7, 74]. In addition, each baseline is equipped with a standard RPF search module [76, 82, 83] that re-finds the target after loss by navigating to the Kalman-filter-predicted position.
Given the prominence of the MPC-based planner, we further reproduce two of its representative variants: (i) MPC w/ trajectory prediction (MPC w/ Traj.) [7], which guides the robot along a predicted target trajectory inferred from historical motion. This planner halts when the predicted trajectory exhibits high uncertainty, thereby prioritizing safety; and (ii) MPC with dynamic search fields (MPC w/ DS.) [48], which employs a dynamic search field to intelligently re-find the target after occlusion—either by overtaking occluders or by temporarily following them until an overtaking opportunity arises.
@@
The current release of Follow-Bench focuses on these map-independent local planners, evaluating their fundamental ability to avoid collisions and stably follow the target across diverse environments. To decouple motion-planning evaluation from perception, we assume perfect knowledge of the target and pedestrian positions whenever visible, with occlusion handled by the simulator’s detection logic. Static obstacles are detected using a clustering-based method applied to 2D LiDAR input, following [80].
@@ III-D Evaluation Protocol and Metrics
Using the re-implemented RPF planners described above, we evaluate their performance across the benchmark scenarios introduced in Sec. III-B. The simulation timestep is fixed at 0.1 s. For each scenario, we first determine the minimum number of simulation steps, , required for the target person to complete their trajectory in the absence of the robot, accounting for natural interactions with other pedestrians. Each scenario is then executed with the robot included for a total of steps, ensuring that the target has sufficient time to complete the trajectory even if temporary interruptions occur.
For each scenario, we conduct 100 independent trials. Across trials, the trajectories of surrounding pedestrians and the placement of static obstacles (when applicable) are randomized to introduce behavioral and environmental diversity. The robot is always initialized at its preferred following position relative to the target. During each trial, we compute the evaluation metrics summarized in Table II. A search-success event is recorded when the robot successfully re-identifies the target within steps after each loss of visual contact. A successful trial is defined as one in which no collisions occur and the robot resumes tracking of the target within steps before the simulation terminates at steps.
@@ IV Analysis of RPF planners
Based on the proposed benchmark, we conduct a series of experiments to evaluate the performance of re-implemented RPF planners. The experiments are designed to analyze how these planners perform under all scenarios (Sec. IV-A), different person-following distances and angles, human densities and environmental occupancy conditions (Sec. IV-B). Here, we simulate human-object and human-human interaction based on ORCA. For simplicity, we show part of the evaluation metrics here, detailed metric evaluation can be found in the released code. In addition, Appendix includes complete evaluation results of all experiments in each scenario.
@@ V-B1 More Context-rich Environments
Currently, Follow-Bench encompasses the most representative scenarios in RPF research, including various target trajectory patterns, crowd dynamics, and environmental layouts. Nevertheless, it primarily evaluates basic motion-planning capabilities in terms of safety and comfort. Future extensions could introduce more context-rich environments, such as hospitals, factories, museums, households, and offices. In these settings, the target person may exhibit application-specific behaviors and interactions with the environment (e.g., visiting a patient, collaborating with colleagues, or shopping), while the robot could be assigned complementary task-oriented goals during following (e.g., delivering medicine, carrying luggage, or guiding to a destination). Such scenarios would demand more sophisticated reasoning and decision-making from RPF planners.
Furthermore, the current benchmark is implemented in a 2D simulator to facilitate efficient evaluation and development of RPF planners due to its simplicity and CPU-only compatibility. It is worth noting that the interaction mechanisms in many high-fidelity simulators [55, 18] remain fundamentally 2D-based—they first generate socially aware 2D human trajectories and subsequently render them with 3D animations for enhanced visualization. Nevertheless, adopting a more realistic 3D simulator (e.g., IsaacSim [98]) would enable richer evaluations of RPF systems under perceptual challenges such as human-trajectory prediction, visual occlusion [48], visual localization on quadruped platforms [38], and semantic understanding [33].
@@ VI Conclusions
Robot person following (RPF) is a fundamental capability in human-robot interaction, with applications spanning personal assistance, security, service, and industrial robotics. This paper provides a comprehensive and structured survey of existing RPF research, with a particular emphasis on two widely recognized and critical criteria: safety and comfort. Guided by these objectives, we review representative scenarios, evaluation metrics, and motion planners. To enable systematic benchmarking of reviewed planners, we introduce Follow-Bench, a unified benchmark implemented in a lightweight simulator that models socially aware interactions, including diverse target trajectory patterns, crowd dynamics, and environmental layouts. To the best of our knowledge, this is the first benchmark to explicitly highlight two essential requirements for RPF: (i) continuously and naturally following a target person, and (ii) proactively avoiding obstacles and pedestrians in socially interactive environments. Follow-Bench allows researchers to develop and evaluate new RPF planners, with flexible support for adding maps, obstacles, agents, metrics, and following strategies. All code and deployment scripts are publicly released.